math question(11 points) A trapezoidal plate, with 9m and 5m bases and 3m height, is submerged vertically in a tank of water so that its t
Answer #1

. Giver am In order to find the area of trapezoid in which the width is increasing from ..sm to am, the width (a) of gle needArea of a Traperoid, A = (base + 2(width)).sx Therefore, . Hydrostatic force hosts 20 E = ? Spg (2-2) da = egļ (2) da - 1000x